Ecola Company uses a job order costing system. Manufacturing overhead is applied on the basis of direct labor cost. Total manufacturing overhead was estimated to be $125,460 for the year; direct labor was estimated to total $153,000.

(1/1) (12/31)
Raw Materials Inventory $ 13,200 $ 10,200
Work in Process Inventory $ 29,200 $ 22,200
Finished Goods Inventory $ 41,200 $ 32,200


The following transactions have occurred during the year.

Raw materials purchases $ 132,000
Direct materials used $ 71,400
Direct labor $ 137,000
Indirect materials used $ 16,200
Indirect labor $ 19,200
Factory equipment depreciation $ 28,200
Factory rent $ 15,200
Factory utilities $ 11,700
Other factory costs $ 8,700

(a) Calculate the predetermined overhead rate

Predetermined Overhead Rate    %

(b) Calculate cost of goods manufactured.

Cost of Goods Manufactured

c) Calculate the over- or underapplied overhead. (Input the amount as positive value.)

  


(d) Calculate adjusted cost of goods sold

Adjusted Cost of Goods Sold   
Answer #1
(a) Predetermined overhead rate=Estimated manufacturing overhead/Estimated direct labor cost=125460/153000=$ 0.82 per direct labor cost
(b) Cost of goods manufactured:
$
Direct materials used 71400
Direct labor 137000
Manufacturing overhead applied (Note:1) 112340
Total manufacturing cost 320740
Add: Work In process inventory on 01/01. 29200
Total manufacturing cost to account for 349940
Less: Work In process inventory on 12/31. 22200
Cost of goods manufactured 327740
Note:1
Manufacturing overhead applied=Direct labor cost*Predetermined overhead rate=137000*0.82=$ 112340
c) Manufacturing overhead applied=$ 112340
Actual manufacturing overhead:
$
Indirect materials used 16200
Indirect labor 19200
Factory equipment depreciation 28200
Factory rent 15200
Factory utilities 11700
Other factory costs 8700
Total 99200
Manufacturing overhead applied > Actual manufacturing overhead
Hence,Overhead is said to be over-applied
Over-applied overhead=112340-99200=$ 13140
d) Adjusted cost of goods sold:
$
Cost of goods manufactured 327740
Add: Finished goods inventory on 01/01. 41200
368940
Less: Finished goods inventory on 12/31. 32200
Unadjusted cost of goods sold 336740
Less: Over-applied overhead 13140
Adjusted cost of goods sold 323600